OpenStudy (kanwal32):

Indicate the correct alternatives for the following-: (listing the question and option)

OpenStudy (kanwal32):

\[for 0<\phi<\pi/2\]

OpenStudy (kanwal32):

if \[x=\sum_{n=0}^{\infty}\cos^{2n} \phi,y=\sum_{n=0}^{\infty}\sin^{2n}\phi \]

OpenStudy (kanwal32):

\[z=\sum_{n=0}^{\infty}\cos ^{2n}\phi(\sin^{2n}\phi)\]

OpenStudy (kanwal32):

then-:(more then one correct) xyz=xz+y xyz=xy+z xyz=x+y+z xyz=yz+x

OpenStudy (kanwal32):

@ganeshie8 @wio @ParthKohli pls help

OpenStudy (kanwal32):

@Luigi0210

OpenStudy (kanwal32):

@ganeshie8 help

OpenStudy (ikram002p):

by 2n , do you mean power or derivative ?

OpenStudy (kainui):

My first idea is to convert it into exponential form and see what happens.

OpenStudy (ikram002p):

@Kainui still not sure 2n is power or n derevative ?

OpenStudy (kainui):

I think if it was derivatives then it wouldn't converge.

OpenStudy (ikram002p):

it would if it was double derevative then for sin note that sin x=- sin '' x cos x = cos '' x ect

OpenStudy (kainui):

Yeah but those sums look like: 1-1+1-1+1-1+... and won't converge

OpenStudy (anonymous):

Answer my messages

OpenStudy (kainui):

Actually if you look at this like a geometric series, you can solve this fairly easily it would seem. For instance, \[\Large x=\frac{1}{1-\cos^2(x)}\] The rest should follow from trig identities fairly nicely.

ganeshie8 (ganeshie8):

\[\large xyz = \dfrac{1}{\sin^2\phi \cos^2\phi \left(1-\sin^2\phi \cos^2\phi\right)}\] ?

ganeshie8 (ganeshie8):

nice :)

OpenStudy (hari5719):

@taylor12344 answer my messages toooo

OpenStudy (kanwal32):

does any1 now about Vn method shortcut

ganeshie8 (ganeshie8):

above is a shortcut method ^

OpenStudy (kanwal32):

yes i got the question

ganeshie8 (ganeshie8):

you just need to know the infinite sum of converging geometric series formula : \(\dfrac{1}{1-r}\)

OpenStudy (kanwal32):

i was able to solve this question on my own

OpenStudy (kanwal32):

but i forgot to close the question

ganeshie8 (ganeshie8):

what method did u use

OpenStudy (kanwal32):

i used sum of infinity in a/1-r and then i started substiuting the value of

OpenStudy (kanwal32):

{sorry} i used the value of phi

OpenStudy (kanwal32):

and ticked the suitble option

ganeshie8 (ganeshie8):

great, we all are on the same page :)

OpenStudy (kanwal32):

ok

OpenStudy (kainui):

It's really not a hard formula to derive. In fact, it's probably one of the most satisfying things in math to do. \[\Large S=1+a+a^2+a^3+... \\ \Large S-1=a+a^2+a^3+... \\ \Large \frac{S-1}{a}= 1+a+a^2+a^3+... \\ \Large \frac{S-1}{a}=S\] Now you can just solve for your sum S with a little algebra. =P

OpenStudy (kanwal32):

yes i solved the question